In a regression analysis, if the predicted value of y is 5 and the corresponding observed value of y is 7, the residual (error) equal to 2.

TRUE OR FALSE ?
Answer #1

True.

The resudial(e) is the difference between the predicted value (ŷ) and the observed value.

Residual = Observed value – predicted value
e = y – ŷ = 7 - 5 = 2